find three consecutive odd integers such that four times the sum of the first and third integer is five more than the product of three and the second interger
Answer by xcentaur(357)

So our three consecutive odd integers will be:
y,(y+2),(y+4)


Acc to the question,
four times the sum of the first and third integer is five more than the product of three and the second integer
4(y+(y+4))= 5 + 3(y+2)
4(2y+4) = 5 + 3y + 6
8y + 16 = 3y + 11
8y - 3y = 11-16
5y = -5
y=-1


Now we have our three numbers as:
-1,1,3


Hope this helps,
Prabhat


[Edit: You can also put (2x+1) instead of y and solve it for x to make sure that you get an odd integer,but its unneccessary]